# An Empirical Study of Testing Practices in Open Source AI Agent Frameworks and Agentic Applications

### v2

@@ 8 Implications
Our empirical study reveals a rapidly evolving testing landscape for FM-based AI agent eco-systems. While practitioners are adapting established software engineering principles, significant gaps and strategic misalignments are evident. The findings from our investigation into testing patterns (RQ1) and component-level focus (RQ2) carry profound implications for the key stakeholders in this ecosystem. In this section, we discuss these implications for practitioners and researchers, outlining actionable pathways to foster a more mature, reliable, and secure agentic ecosystem.
@@ 8.2 Implications for Researchers
Researchers should conduct empirical studies to diagnose the barriers preventing the adoption of novel, agent-specific testing patterns. The contrast between the potential of patterns like DeepEval and Hyperparameter Control and their near-zero adoption (Section 6) points to a significant gap between the state-of-the-art and the state-of-the-practice. Is this gap caused by a lack of awareness, tooling immaturity, perceived complexity, or a cultural adherence to traditional testing paradigms? Future research, employing qualitative methods like surveys and interviews with practitioners, could uncover these root causes. The resulting insights would be invaluable for designing more intuitive testing tools, developing targeted educational materials, and creating theories of technology adoption tailored to the rapidly evolving FM-based AI agent engineering landscape.
Researchers should formalize a comprehensive testing methodology for agentic systems that accounts for their unique architectural and behavioral characteristics. Our study revealed two foundational shifts in testing practice: the strategic adaptation of “less strict” patterns (RQ1) and the inversion of testing effort away from the non-deterministic model and towards the deterministic infrastructure (RQ2). These emergent practices currently lack a unifying theoretical framework. Researchers are positioned to develop a “Theory of Agent Testing” that formalizes these observations. Such a theory would provide prescriptive guidance on which testing patterns are most effective for which canonical components. For instance, using Test Double for Boundary Artifacts, Negative Testing for Constitutive Entities, and DeepEval for Observable Properties. By creating a systematic, evidence-based methodology, researchers can move the field from ad-hoc practices to a disciplined engineering approach. Though we have seen some early research about this topic (Shamim and Singhal, 2024), but there are a lot of work needed.
@@ A.1.1 Hyperparameter Control
5 ceo = Agent(name=’CEO’, tools=tools, instructions="You are an agent that tests concurrent API calls. You must say ’success’ if the output contains headers, and ’error’ if it does not and **nothing else**.")
@@
7 result = agency.get_completion("Please call PrintHeaders tool TWICE at the same time in a single message. If any of the function outputs do not contains headers, please say ’error’.")
8 self.assertTrue(result.lower().count(’error’) == 0, agency.main_thread.thread_url)
@@ A.2.2 DeepEval
Listing 5: DeepEval: Using DeepEval for verifying the company name name in a job description by a resume retriever agent.
@@ 11 GEval(
19 f"Check that the output contains the company name ’{company}’",

### v3

@@ 8 Implications
Our empirical study reveals a rapidly evolving testing landscape for FM-based AI agent ecosystems. While practitioners are adapting established software engineering principles, significant gaps and strategic misalignments are evident. The findings from our investigation into testing patterns (RQ1) and component-level focus (RQ2) carry profound implications for the key stakeholders in this ecosystem. In this section, we discuss these implications for practitioners and researchers, outlining actionable pathways to foster a more mature, reliable, and secure agentic ecosystem.
@@ 8.1.1 For Agent Framework Developers
However, the reluctance to adopt these patterns can be understood in light of practical constraints. The underlying use of LLM-as-a-Judge in these patterns introduces substantial overhead. Unlike traditional unit tests, which execute in milliseconds, semantic evaluation requires invoking a secondary LLM to judge the output produced by the first LLM. This creates a “token tax,” i.e., the judge requires the full context plus the generated output, effectively doubling the number of tokens consumed and making such setups impractical for large-scale or production-grade testing (Rajbahadur et al., 2024). When reasoning models are used as LLM-as-a-Judge, costs can escalate drastically; recent benchmarks indicate that using top-tier models (e.g., GPT-5-reasoning) for evaluation can cost nearly USD-$79.00$ per 1,000 evaluations (Huang et al., 2025). Consequently, latency can become a bottleneck as multiple inference cycles might be required to complete the testing (Huyen, 2024).
Beyond resource constraints, reliability remains an issue. AI judges are also probabilistic and prone to criteria ambiguity along with specific biases, e.g., verbosity bias, ordering or self-preference, and often require a model stronger than the application itself to provide a valid critique (Huyen, 2025a). Consequently, these judges still require periodic human oversight to ensure the evaluator does not drift (Huyen, 2025b).
To address the “token tax” and high inference latency, frameworks can facilitate the use of smaller, specialized judge models using curriculum engineering (Rajbahadur et al., 2024). Several teams in practice have used specialized, smaller, and less expensive models to effectively validate the outputs of stronger, more expensive models without compromising validity (Huyen, 2024). Additionally, frameworks can support “spot-checking” capabilities by leveraging established test prioritization techniques and hierarchical test trees (Yoo and Harman, 2012; Cruciani et al., 2019). This may allow developers to evaluate only a statistically significant subset of responses rather than the entire test suite, achieving a pragmatic balance between cost and confidence. Furthermore, to mitigate latency bottlenecks, testing tools should support asynchronous architectures, e.g., offloading expensive semantic evaluations to nightly runs or background processes rather than blocking immediate deployment pipelines. Finally, to enhance reliability, frameworks can explicitly manage evaluator bias by adopting strategies such as automatic input permutation, prompt management, and comparative evaluation. This will shift the complexity of robust, cost-effective testing from application developers to the framework itself.
@@
Framework developers should mitigate the gap between testing of deterministic components and non-deterministic agent artifacts. In our study, we observed that non-deterministic FMs receive comparatively less attention in testing effort. Beyond tooling limitations, this pattern may be a consequence of deeper cultural and cognitive factors. Prior studies show that practitioners prioritize the controllability and observability of the subject under test when deciding which modules to test (Garousi et al., 2019). Controllability ensures consistent outputs for controlled inputs, while observability enables straightforward inspection and interpretation of test results. Deterministic components naturally satisfy these criteria, whereas non-deterministic FMs violate long-standing testing norms rooted in reproducibility and oracle-based validation. In this sense, the shift of testing effort can be understood as a strategic response to current capability limitations rather than a contradiction of them, as prior work also shows that testing strategies often emerge from practical constraints (Whyte and Mulder, 2011). Because non-deterministic components are harder to control and observe, developers seem to adapt by prioritizing artifacts whose behavior can be verified more feasibly under existing testing constraints.
This mismatch is further reinforced by developers’ implicit acceptance of LLM non-determinism as an inherent system property rather than a testable fault. Recent surveys indicate that practitioners struggle to integrate agents into CI/CD pipelines precisely because these pipelines presuppose deterministic behavior (Pan et al., 2025). As a result, developers often adopt compensatory practices, such as limiting the number of reasoning steps or introducing human-in-the-loop validation, rather than attempting to address or test the non-determinism (Pan et al., 2025). Further, qualitative evidence from open-source development environments reinforces this interpretation and illustrates how practitioners attempt to address these challenges in practice. For example, framework developers explicitly discuss “flaky” behavior arising from non-determinism and attempt to mitigate it  [^5] https://github.com/langchain-ai/langgraph/pull/319, often by mocking LLM clients to enforce consistent reproducibility during testing  [^6] https://github.com/microsoft/autogen/pull/4096. Similarly, practitioners discuss the difficulty of verifying whether agents actually executed tools or merely hallucinated results  [^7] https://github.com/crewAIInc/crewAI/pull/3388. Together, these observations suggest that the lower testing effort devoted to non-deterministic components is not solely a consequence of technical limitations. It also reflects a long-standing culture among developers of expecting deterministic behavior in tests, as well as the implicit acceptance of LLM non-determinism as a given system property. This combination highlights the need for framework developers to actively mitigate the gap between testing of deterministic components and non-deterministic agent artifacts.
@@ 8.2 Implications for Researchers
Researchers should conduct empirical studies to diagnose the barriers preventing the adoption of novel, non-determinism-specific testing patterns. The contrast between the potential of patterns like DeepEval and Hyperparameter Control and their near-zero adoption (Section 6) points to a significant gap between the state-of-the-art and the state-of-the-practice. Is this gap caused by a lack of awareness, tooling immaturity, perceived complexity, or a cultural adherence to traditional testing paradigms? Future research, employing qualitative methods like surveys and interviews with practitioners, could uncover these root causes. The resulting insights would be invaluable for designing more intuitive testing tools, developing targeted educational materials, and creating theories of technology adoption tailored to the rapidly evolving FM-based AI agent engineering landscape.
Researchers should formalize a comprehensive testing methodology for agentic systems that accounts for their unique architectural and behavioral characteristics. Our study revealed two foundational shifts in testing practice: the strategic adaptation of “less strict” patterns (RQ1) and the inversion of testing effort away from the non-deterministic model and towards the deterministic artifacts and infrastructure (RQ2). To enhance the implications of our findings, we synthesized best practices from traditional Software Engineering (SE), classical Machine Learning (ML), and Large Language Model (LLM) application domains. We queried Google Scholar using terms such as “unit testing best practices”, “machine learning system testing best practices”, “testing llm applications”, and “testing large language model based software”. Table 9 maps these recommendations to the adoption behaviors observed in agent frameworks and agentic applications in our study. Our empirical results show that agent frameworks and agentic applications fully adopt, partially adopt, adopt in a simplified form, or do not adopt these inherited practices.
From traditional SE, we observe partial adoption of deterministic testing, code coverage, and systematic test design in both RQ1 and RQ2. These practices are applied primarily to deterministic system shell components, but are not used across full agent workflows. The strict recommendation of deterministic oracles from SE (Daka and Fraser, 2014) is often infeasible for non-deterministic model components in FM-based agents.
From classical ML recommendations, regression testing practices are partially adopted in agent systems, whereas metamorphic and differential testing are not, as property-based validation remains difficult to define for open-ended agent outputs (Chandrasekaran et al., 2023). Negative testing is partially adopted, particularly for constitutive entities, whereas adversarial fuzzing remains rare in practice.
From the LLM application domain, several recommendations are either fully or partially adopted. For instance, probabilistic assertions on top of deterministic tests are being partially adopted in agent frameworks and agentic applications. For example, in Section A.2.2, we illustrate a real-world test function that adopts probabilistic testing, which represents a paradigm shift over deterministic testing. Also, layered testing is fully adopted across most systems. Curriculum-based AI judges and controlled execution for repeatability are adopted in simplified form, for example, through patterns such as DeepEval, Parameterized Testing, and Mock Assertion. In contrast, aggregated oracles and prompt perturbation for robustness have not yet been adopted. Component-aware testing strategies are adopted in simplified form, often implicitly, without formal guidance.
Overall, this comparison suggests that agent testing is best understood as a hybrid regime. It adopts mature deterministic testing practices for infrastructure, adapts evaluation and oracle strategies from LLM testing to accommodate expected non-determinism, and finds some classical ML testing strategies not directly transferable without component-specific reformulation.
Table 9: Comparison of academic testing recommendations and their adoption in agentic systems. The bracketed numbers in the Domain column denote the following sources: Traditional SE: [1] (Daka and Fraser, 2014); Classical ML: [2] (Chandrasekaran et al., 2023), [3] (Openja et al., 2024); LLM Applications: [4] (Dobslaw et al., 2025), [5] (Ma et al., 2025), [6] (Rajbahadur et al., 2024).
Recommendation Domain Adoption Status Remarks
Use deterministic tests SE [1] Adopted Partially Strictly applied to deterministic infrastructure (System Shell), but infeasible for the non-deterministic model core.
Code coverage and systematic test design SE [1] Adopted Partially High coverage is pursued for Resource Artifacts, but widely ignored for Triggers
Metamorphic and differential testing ML [2] Not Adopted Formal property-based relations are difficult to define for open-ended agent outputs.
Regression testing practices ML [3] Adopted Partially Observed in deterministic artifacts, though often missing critical regression selection strategies for prompts.
Negative and adversarial testing ML [2] Adopted Partially Negative testing is common for constitutive entity, but adversarial fuzzing is rare.
Aggregated oracles (voting, variance) LLM [4] Not Adopted Systematic aggregation (e.g., majority voting) is never formalized.
Prompt perturbation for robustness LLM [4] Not Adopted Systematic perturbation to test prompt fragility is theoretically recommended but absent in practice.
Probabilistic assertions LLM [5] Adopted Fully Widely adopted as “less strict” assertions (semantic similarity, embedding checks).
Layered testing (System Shell vs. Model) LLM [5] Adopted Fully Traditional testing patterns for deterministic components (system shell) and less strict patterns for non-deterministic components (model)
Curriculum-based AI judges LLM [6] Adopted in Simplified Form Advanced judge training is not adopted; however, simpler judge implementations (e.g., DeepEval) are emerging.
Controlled execution for repeatability LLM [6] Adopted in Simplified Form Repeatability is approximated through parameterized testing and controlled configuration through test doubles, mock assertion and hyper-parameter control.
Component-aware testing strategies LLM [5] Adopted in Simplified Form Developers implicitly tailor testing to agent components (e.g., boundary artifacts).
We therefore call for a unified “Theory of Agent Testing” that makes these transfer boundaries explicit and provides prescriptive guidance linking canonical agent components to suitable testing strategies, thereby moving practice from ad hoc solutions toward disciplined, evidence-based testing.
@@ 9.1 External Threats
Our study focuses on identifying how different components of agentic systems are tested in practice, rather than empirically evaluating how variations or failures in those components affect system behavior. For example, while we report limited prompt testing in current practice, analyzing how prompt variations influence agent behavior is outside the scope of this study. Instead, we refer to prior component-specific research that investigates such effects.
@@ 9.2 Construct Threats
Finally, our reliance on the test_* filename convention to identify test files constitutes a construct validity threat. While this aligns with standard practices in Python ecosystems (e.g., pytest discovery rules) and prior empirical studies (Jebnoun et al., 2020; Openja et al., 2024), it may exclude repositories that employ other naming schemes or embed tests within production code. As a result, our findings are scoped to repositories that follow explicit prefix-based test file naming conventions. Investigating alternative testing structures and assessing whether repositories that follow other conventions exhibit different testing practices remain important directions for future work.
@@ 10 Conclusion
11 Data Availability Statement
The datasets generated and analyzed during this study are available in the replication package.
@@ A.1.1 Hyperparameter Control
5 ceo = Agent(name=\’CEO\’, tools=tools, instructions="You are an agent that tests concurrent API calls. You must say success if the output contains headers, and error if it does not and **nothing else**.")
@@
7 result = agency.get_completion("Please call PrintHeaders tool TWICE at the same time in a single message. If any of the function outputs do not contains headers, please say error’.")
8 self.assertTrue(result.lower().count(’error’) == 0, agency.main_thread.thread_url)’
@@ A.2.2 DeepEval
Listing 5: DeepEval: Example of probabilistic assertion using DeepEval to verify that the retrieved output contains the expected company name in a resume retrieval agent.
@@ 11 GEval(
19 f"Check that the output contains the company name ‘company’"
20 f"Check that the output does not contain any company name other than ’company’"
@@ A.2.7 Value Range Analysis
Appendix B Canonical Agent Architecture
B.1 Belief Base
B.2 Goal
B.3 Plan
B.4 Trigger
B.5 Context
B.6 Plan Body
B.7 Internal Action
B.8 External Action
B.9 Communicative Action
B.10 Resource Artifact
B.11 Coordination Artifact
B.12 Boundary Artifact
B.13 Observable Property & Signal
B.14 Registry